# SimSort: A Powerful Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

### v1

@@ A.3 Spike Identification
A.4 Definitions of Metrics
To evaluate the performance of spike sorting algorithms, we adopted standard metrics as defined in the context of spike sorting and ground-truth comparisons (Magland et al., 2020; Buccino et al., 2020). These metrics included accuracy, precision, and recall, which were derived based on the matching between the sorted spike train and the ground-truth spike train.
@@ Appendix B Supplementary Figures
Figure S1: (a) Accuracy of spike sorting on the Hybrid dataset. SimSort outperforms HerdingSpikes2 (p = 0.00), JRClust (p = 0.02), and KiloSort2 (p = 0.00) in static recordings, with no significant difference compared to IronClust (p = 0.45), KiloSort (p = 0.71), MountainSort4 (p = 0.48), SpykingCircus (p = 0.27), and Tridesclous (p = 0.12). In the drift condition, SimSort significantly outperforms HerdingSpikes2 (p = 0.00), JRClust (p = 0.00), KiloSort2 (p = 0.00), MountainSort4 (p = 0.00), and Tridesclous (p = 0.00), with no significant difference compared to IronClust (p = 0.69), KiloSort (p = 0.24), and SpykingCircus (p = 0.06). (b) Accuracy of spike sorting on the WaveClus dataset. SimSort shows no significant improvement in the easy subset compared to KiloSort (p = 0.14), KiloSort2 (p = 0.22), MountainSort4 (p = 0.86), and MountainSort5 (p = 0.69). In the difficult subset, SimSort significantly outperforms KiloSort (p = 0.00) and KiloSort2 (p = 0.00), with no significant difference compared to MountainSort4 (p = 0.22) and MountainSort5 (p = 0.13). Statistical analysis was performed using an independent two-tailed $t$-test.

### v2

@@ A.4 Spike Identification
A.5 Definitions of Metrics
To evaluate the performance of spike sorting algorithms, we adopted standard metrics as defined in the context of spike sorting and ground-truth comparisons [41, 44]. These metrics included accuracy, precision, and recall, which were derived based on the matching between the sorted spike train and the ground-truth spike train.
@@ A.6 Hyperparameter Search
A.7 Infrastructures
Our simulation dataset are generated using CPU clusters for approximately 5000 CPU hours, which can be parallelized. Our models are trained on NVIDIA A100/V100 GPU, approximately 20 GPU hours for each time.
@@ Appendix B Supplementary Figures
Figure S1: (a) Accuracy of spike sorting on the Hybrid dataset. SimSort significantly outperforms HerdingSpikes2 ($p=0.0000^{*}$), JRClust ($p=0.0000^{*}$), KiloSort2 ($p=0.0000^{*}$), and Tridesclous ($p=0.0001^{*}$) in static recordings, and also shows a significant advantage over IronClust ($p=0.0296^{*}$). There is no significant difference compared to KiloSort ($p=0.5023$), MountainSort4 ($p=0.1799$), and SpykingCircus ($p=0.1356$). In the drift condition, SimSort significantly outperforms HerdingSpikes2 ($p=0.0000^{*}$), JRClust ($p=0.0000^{*}$), KiloSort ($p=0.0285^{*}$), KiloSort2 ($p=0.0000^{*}$), MountainSort4 ($p=0.0004^{*}$), SpykingCircus ($p=0.0111^{*}$), and Tridesclous ($p=0.0007^{*}$), with no significant difference compared to IronClust ($p=0.3398$). (b) Accuracy of spike sorting on the WaveClus dataset. In the easy subset, SimSort shows no significant improvement compared to KiloSort ($p=0.0770$), KiloSort2 ($p=0.1569$), MountainSort4 ($p=0.6137$), and MountainSort5 ($p=0.4424$). In the difficult subset, SimSort significantly outperforms KiloSort ($p=0.0007^{*}$), KiloSort2 ($p=0.0004^{*}$), and MountainSort5 ($p=0.0449^{*}$), with no significant difference compared to MountainSort4 ($p=0.0935$). Statistical analysis was performed using a paired two-tailed $t$-test.
@@
Figure S4: Impact of dimensionality reduction and clustering algorithms on the hybrid dataset. MS denotes Mean Shift, SC denotes Spectral Clustering. Comparing different combinations of dimensionality reduction methods (None, PCA, $t$-SNE, UMAP) and clustering algorithms (GMM, KMeans, Mean Shift, and Spectral Clustering) on static (left) and drift (right) subset. Here, “None” indicates that raw features learned by SimSort were used directly without dimensionality reduction.
@@ Appendix C More Showcases
Figure S17: Visualization of spike sorting performance on real recording data.